OpenStudy (anonymous):

g(f(h(-3))) f(x)= 2x g(x)= -3x+4 h(x)= 2x-8

OpenStudy (lieutenantgeneral):

f(x)= 2x = real number odd g(x)= -3x+4 = 4 h(x)= 2x-8 = 2x = h(x) + 8 h(x) = 2 (x - 4)

OpenStudy (anonymous):

hahaha im so confused.. so how do I find the answer? I don't understand this lesson at all..

OpenStudy (lieutenantgeneral):

www.wolframalpha.com lol

OpenStudy (anonymous):

h(x)= 2x-8 = 2x = h(x) + 8 h(x) = 2 (x - 4) should I type this in? lol